### v1

# Bridging Relevance and Reasoning: Rationale Distillation in Retrieval-Augmented Generation

Abstract
The reranker and generator are two critical components in the Retrieval-Augmented Generation (i.e., RAG) pipeline, responsible for ranking relevant documents and generating responses. However, due to differences in pre-training data and objectives, there is an inevitable gap between the documents ranked as relevant by the reranker and those required by the generator to support answering the query. To address this gap, we propose RADIO, a novel and practical preference alignment framework with RAtionale DIstillatiOn. Specifically, We first propose a rationale extraction method that leverages the reasoning capabilities of Large Language Models (LLMs) to extract the rationales necessary for answering the query. Subsequently, a rationale-based alignment process is designed to rerank the documents based on the extracted rationales, and fine-tune the reranker to align the preferences. We conduct extensive experiments on two tasks across three datasets to demonstrate the effectiveness of our approach compared to baseline methods. Our code is released online  [^1] The link will be released after publication to ease reproduction.
Bridging Relevance and Reasoning: Rationale Distillation in Retrieval-Augmented Generation
Pengyue Jia1111equal contribution, Derong Xu1111equal contribution, Xiaopeng Li1111equal contribution, Zhaocheng Du2, Xiangyang Li2, Xiangyu Zhao1, Yichao Wang2, Yuhao Wang1, Huifeng Guo2, Ruiming Tang2, 1City University of Hong Kong, 2Huawei Noah’s Ark Lab
1 Introduction
Large Language Models (LLMs), pretrained on massive datasets, have demonstrated exceptional reasoning and text generation capabilities, as evidenced by prior research Zhao et al. (2023). These models also adhere to the scaling laws, exhibiting improvements in performance and intelligence as the number of parameters increases Kaplan et al. (2020). Retrieval-augmented generation (RAG) builds upon these capabilities by integrating information retrieval mechanisms with generative models, such as LLMs. This approach not only mitigates the problem of hallucination in text generation but also enhances the system’s adaptability to dynamically evolving information needs, making it a robust solution for tasks requiring both accuracy and contextual relevance Gao et al. (2023).
However, the RAG pipeline typically assembles components (e.g., the reranker and generator Fan et al. (2024)) that have been pretrained separately. Due to differences in their pretraining data and optimization objectives, these components often exhibit varying preferences within the pipeline, which can impact the overall effectiveness of the system. Specifically, pretrained rerankers Xiao et al. (2023) excel at evaluating the relevance between queries and documents. However, the documents identified as “relevant” under this criterion may not provide the necessary support for reasoning to derive an accurate answer to the query. Bridging this gap between the reranker’s relevance-based rankings and the generator’s requirements for reasoning support presents a significant challenge that must be addressed to improve the RAG pipeline’s performance. Recent studies try to address this gap by training a bridge model Ke et al. (2024) or using LLM-based Zhang et al. (2024) and retrieval-based scores to fine-tune RAG components Dong et al. (2024). Additionally, while some other methods are not specifically designed to tackle this problem, they can indirectly help bridge the gap. These approaches can use response quality Ma et al. (2023) or perplexity distillation Izacard et al. (2023); Shi et al. (2023) as signals to fine-tune the reranker. While these approaches show promise, they face critical limitations: relying solely on the connection between the query or answer and the document to determine the document’s importance captures only surface-level relationships, neglecting the deeper, rationale-based reasoning.
To this end, we propose RADIO, a novel and practical preference alignment framework with rationale distillation in retrieval-augmented generation. First, to efficiently extract the rationales necessary for answering a query, we leverage the query and its ground truth answer as contextual information. Then, we generate the rationale required to derive the correct answer with LLMs. Second, to mitigate the preference misalignment between the reranker and generator while ensuring the solution remains practical, we rerank the documents based on the extracted rationale and fine-tune the reranker using this rationale-based ranking. This step distillates rationales from generators to rerankers, and aligns the reranker with the generator’s information needs for answering the query effectively.
RADIO effectively addresses the preference inconsistency between RAG components by first generating a comprehensive rationale and then fine-tuning the reranker based on the extracted rationale. This approach considers the deeper reasoning behind answers. We evaluate RADIO on two tasks across three datasets: Open-domain QA (NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017)) and Multi-choice questions (MMLU Hendrycks et al. (2020)). The results validate the superiority of our method compared to other state-of-the-art baselines. Our contributions can be summarized as follows:
  • We propose RADIO, a novel and practical framework designed to address the preference misalignment of different components in RAG pipelines.
  • We introduce rationale distillation within the RAG framework, which is an effective approach that leverages explicit textual rationales as signals to align the preferences of different components in RAG.
  • Extensive experiments are conducted on two tasks across three datasets to demonstrate the effectiveness and transferability of RADIO.
2 Methodology
In this section, we introduce rationale distillation from generators to rerankers to mitigate the preference misalignment between rerankers and generators in RAG. Specifically, we first demonstrate the task definition of RAG in Section 2.1. Then we give an overview of our proposed framework in Section 2.2, introduce the rationale extraction method in Section 2.3, and detail the rationale-based alignment in Section 2.4 and optimization in Section 2.5.
2.1 Task Definition
To address the hallucination problem and enhance adaptability to dynamic information of LLMs, RAG systems have been proposed. These systems enhance generative models by introducing additional contextual information retrieved based on a given query $q$. Specifically, when a query $q$ is input into the RAG pipeline, the retriever $\mathcal{R}_{\text{retriever}}$ first identifies relevant documents by calculating similarity scores and top-$k_{1}$ selection. The process can be formalized as follows:
$\mathcal{D}_{\text{retriever}}=\{\boldsymbol{d}_{i}\mid\boldsymbol{d}_{i}\in\text{Top-}k_{1}(\text{score}_{\text{retriever}}(q,d))\}$ (1)
where $q$ and $d$ are the query and document, $\text{score}_{\text{retriever}}$ denotes the score function in retriever, $\boldsymbol{d}_{i}$ means the $i$-th document in corpus and $\mathcal{D}_{\text{retriever}}$ is the documents set output by retriever $\mathcal{R}_{\text{retriever}}$.
To eliminate contextually irrelevant noise and provide more precise contextual information, the initially filtered documents will be further reranked by the reranker:
$\mathcal{D}_{\text{reranker}}=\{\boldsymbol{d}_{j}\mid\boldsymbol{d}_{j}\in\text{Top-}k_{2}(\text{score}_{\text{reranker}}(q,d))\}$ (2)
where $\mathcal{D}_{\text{reranker}}$ is the documents selected by reranker, $\text{score}_{\text{reranker}}$ denotes the score function in reranker. Note that the $j$th document $\boldsymbol{d}_{j}$ is in $\mathcal{D}_{\text{retriever}}$ (i.e., $d_{j}\in\mathcal{D}_{\text{retriever}}$) and $k_{2}$ is the number of documents selected by reranker, which is smaller than $k_{1}$ used by the retriever.
Finally, the documents filtered by the reranker, along with the original query, will be fed into the generator as contextual information to help generate the final response:
$\hat{y}=\mathcal{G}(q,\mathcal{D}_{\text{reranker}})$ (3)
where $\hat{y}$ is the generated response and $\mathcal{G}$ denotes the generator.
It is worth noting that the documents selected by the reranker directly influence the generator’s input. Therefore, in this work, we aim to align the preferences of the reranker and generator to enhance their consistency. This alignment improves the overall accuracy of the RAG system’s responses.
2.2 Framework Overview
Figure 1: Overview of RADIO.
The overview of RADIO is depicted in Figure 1. RADIO is consisted of two phases: rationale extraction (Figure 1(a)) and rationale-based alignment (Figure 1(b)).
In the rationale extraction process, we aim to obtain the correct rationale required to answer the query accurately. To achieve this, we combine the query with the ground truth answer and input them into LLM. By using the correct answer as guiding context in the prompt, we can obtain the rationale with greater precision. This approach leverages the LLM’s reasoning capabilities more effectively, ensuring that the generated rationale closely aligns with the requirements for deriving the correct answer.
In the rationale-based alignment process, our goal is to use the generated rationale to guide the reranker, enabling it to select documents that better support the generator in answering the query. Specifically, we leverage the generated rationale as a signal to rerank documents. The reranked documents will be used to fine-tune the reranker, addressing the preference misalignment between the reranker and the generator. By aligning these components, the process ensures that the selected documents are not only contextually relevant but also optimally supportive for the generator’s reasoning and response generation.
2.3 Rationale Extraction
Rationales are critical components of LLM reasoning processes and have been shown to significantly enhance the accuracy of LLM-generated responses. This perspective is supported by existing works such as Chain-of-Thoughts (CoT Wei et al. (2022)) and O1 Zhong et al. (2024). Existing work Shi et al. (2023); Ma et al. (2023); Dong et al. (2024) has primarily focused on the initial relationships between queries and documents or indirect relationships between answers and documents, while overlooking rationales—a crucial intermediary component in the reasoning process. Motivated by this, we aim to extract rationales as signals to align the preferences of different components in the RAG pipeline.
To accurately extract the rationale necessary for answering a query and deriving the correct answer, we combine the query with the ground truth answer as contextual information of LLMs, as shown in Figure 1(a). The prompt template used for this process is as follows: "You are a professional QA assistant. Given a question and the ground truth answer, you can output the rationale why the ground truth answer is correct. Question: {question}. Answer: {answer}. Rationale: ". By doing so, we effectively bridge the gap between the query and the answer by generating the necessary rationale. This rationale accurately supports the reasoning process required to derive the correct answer from the query.
2.4 Rationale-based Alignment
Given the extracted rationale $r$, a key challenge lies in effectively and efficiently utilizing it to improve preference consistency within the RAG pipeline. In this section, we propose a rationale-based alignment approach, where the rationale serves as a signal to fine-tune the reranker. This enables the reranker to identify and prioritize supportive documents that facilitate the generator in producing accurate responses. Specifically, we first use the retriever $\mathcal{R}_{\text{retriever}}$ to retrieve $k_{1}$ relevant documents based on the query:
$\mathcal{D}_{\text{retriever}}=\mathcal{R}_{\text{retriever}}(q,\mathbf{\mathcal{C})}$ (4)
where $\mathcal{D}_{retriever}$ is the document set retrieved by $\mathcal{R}_{\text{retriever}}$ based on query $q$ and corpus $\mathbf{\mathcal{C}}$, and $|\mathcal{D}_{retriever}|=k_{1}$.
Next, to facilitate the comparison of similarity between different documents and the rationale, we use a text encoder to convert both the documents and the rationale into dense vectors.
$\boldsymbol{e}^{\text{document}}_{i}=\text{Encoder}(\boldsymbol{d}_{i})$ (5)
$\boldsymbol{e}^{\text{rationale}}=\text{Encoder}(r)$ (6)
where $\boldsymbol{e}^{\text{document}}_{i}$ and $\boldsymbol{e}^{\text{rationale}}$ denote the representations of $i$th document and rationale. $\boldsymbol{d}_{i}$ is the $i$th document and $r$ represents the extracted rationale.
Then, we calculate the semantic similarity between each document and the rationale. Here, we use cosine similarity, denoted as $\text{sim}(\cdot)$. The calculated scores indicate the degree to which each document supports generating the correct answer, with higher scores reflecting stronger support. We also linearly interpolate the score of documents with their retrieval score in the retrieval stage by weighted score sum.
$\boldsymbol{s}_{i}^{\text{rationale}}=\text{sim}(\boldsymbol{e}^{\text{document}}_{i},\boldsymbol{e}^{\text{rationale}})$ (7)
$\boldsymbol{s}_{i}^{\text{retriever}}=\text{score}_{\text{retriever}}(q,\boldsymbol{d}_{i})$ (8)
$\boldsymbol{s}_{i}^{\prime}=\alpha\boldsymbol{s}_{i}^{\text{rationale}}+(1-\alpha)\boldsymbol{s}_{i}^{\text{retriever}}$ (9)
where $\boldsymbol{s}_{i}^{\text{rationale}}$, $\boldsymbol{s}_{i}^{\text{retriever}}$, $\boldsymbol{s}_{i}^{\prime}$ represent the rationale similarity score, retrieval score, and final score for $i$th document. $\text{score}_{\text{retriever}}(\cdot)$ is the score function in retriever and $\alpha$ is a hyperparameter used for integration. Note that we apply min-max normalization in this work to both rationale score and retrieval score before integration.
Next, we rerank the documents based on their scores. Following the previous sampling method Top-k shifted by N Moreira et al. (2024b), we select the top-ranked document as the positive sample and then shift by $n$ documents and sample $m$ negative samples from the subsequent documents to construct positive-negative pairs for fine-tuning the reranker. This process can be represented as:
$\boldsymbol{d}_{\text{pos}}=\boldsymbol{d}_{i},\text{where}\ i=\arg\max_{i}\boldsymbol{s}_{i}^{\prime}$ (10)
$\{\boldsymbol{d}_{\text{neg}}\}=\text{Sample}_{m}(\{\boldsymbol{d}_{i}|\text{rank}(\boldsymbol{d}_{i})\textgreater n\})$ (11)
where $\boldsymbol{d}_{\text{pos}}$ and $\boldsymbol{d}_{\text{neg}}$ are the sampling positive and negative documents, $\text{Sample}_{m}(\cdot)$ denotes a sampling operation that selects $m$ negative documents from the set of documents ranked lower than $n$.
2.5 Optimization
Following BGE embedding Xiao et al. (2023) and QA Ranking Benchmark Moreira et al. (2024a), we use InfoNCE as our optimization objectives to fine-tune reranker:
$f(q,d)=\text{exp}(\phi(q,d)/\tau)$ (12)
$L=-\text{log}\frac{f(q,d^{+})}{f(q,d^{+})+\sum_{i=1}^{N}f(q,d_{i}^{-})}$ (13)
where $d^{+}$ and $d^{-}$ represent the positive and negative document, $\tau$ is the temperature parameter, and $N$ denotes the number of negative documents.
3 Experiments
3.1 Datasets and Metrics
To evaluate RADIO with other methods, we conduct experiments on two tasks across three datasets: Open-domain QA (NQ Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017)) and Multi-choice questions (MMLU Hendrycks et al. (2020)). Deatiled dataset descriptions are given in Appendix A.1. Following previous work Ma et al. (2023); Shi et al. (2023), we report EM and F1 scores for Open-domain QA datasets and EM for MMLU.
3.2 Baselines
To verify the effectiveness of RADIO, we conduct experiments with the following baseline methods: Base Xiao et al. (2023), Atlas Izacard et al. (2023), REPLUG Shi et al. (2023), Trainable rewrite-retrieve-read (RRR) Ma et al. (2023), ARL2 Zhang et al. (2024), and DPA-RAG Dong et al. (2024). The detailed introduction of baselines is given in Appendix A.2.
Table 1: Overall experiments.
Method NQ TriviaQA
gte-base gte-large bge-reranker-base gte-base gte-large bge-reranker-base
EM F1 EM F1 EM F1 EM F1 EM F1 EM F1
Base 0.2931 0.4046 0.2798 0.3935 0.3371 0.4603 0.5449 0.6374 0.5495 0.6414 0.6114 0.7120
Atlas 0.3338 0.4587 0.3418 0.4677 0.3521 0.4832 0.5823 0.6752 0.6004 0.6972 0.6083 0.7063
REPLUG 0.3257 0.4484 0.2607 0.3670 0.3427 0.4753 0.5679 0.6578 0.5248 0.6171 0.6032 0.7004
RRR 0.3374 0.4608 0.3299 0.4578 0.3438 0.4754 0.5801 0.6716 0.5358 0.6237 0.6099 0.7091
ARL2 0.3413 0.4688 0.3515 0.4804 0.3568 0.4885 0.6079 0.7086 0.6107 0.7120 0.6137 0.7149
DPA-RAG 0.3391 0.4674 0.3385 0.4710 0.3462 0.4793 0.6080 0.7076 0.6097 0.7119 0.6149 0.7169
RADIO (Ours) 0.3512 0.4790 0.3565 0.4850 0.3665 0.4917 0.6084 0.7095 0.6128 0.7137 0.6154 0.7151
Table 2: Experimental results on MMLU. EM is reported as the metric. The source datasets used to fine-tune rerankers are Open-domain QA datasets NQ and TriviaQA.
Method MMLU (Source Dataset NQ) MMLU (Source Dataset TriviaQA)
Humanities Social STEM Other ALL Humanities Social STEM Other ALL
Base 0.4089 0.6867 0.5147 0.6650 0.5502 0.4089 0.6867 0.5147 0.6650 0.5502
Atlas 0.3985 0.6935 0.5074 0.6563 0.5447 0.3966 0.6822 0.4868 0.654 0.5364
REPLUG 0.4102 0.6854 0.5065 0.6590 0.5473 0.3977 0.6744 0.4821 0.6466 0.5323
RRR 0.4079 0.6913 0.5116 0.6572 0.5484 0.4132 0.6926 0.5005 0.6501 0.5464
ARL2 0.4147 0.7016 0.5106 0.6630 0.5540 0.4012 0.6951 0.5011 0.6639 0.5462
DPA-RAG 0.4157 0.701 0.5078 0.6652 0.5541 0.4189 0.6932 0.5062 0.6746 0.5552
RADIO (Ours) 0.4172 0.7013 0.5080 0.6717 0.5562 0.423 0.6942 0.5090 0.6678 0.5559
3.3 Backbone Rerankers
To validate the generality and adaptability of RADIO, we selected three different rerankers as the backbone models for our experiments: (1) gte-base Li et al. (2023): a reranker model proposed by Alibaba DAMO Academy, with 109M parameters and 768 embedding dimensions. (2) gte-large Li et al. (2023): the larger version of gte-base, with 335M parameters and 1024 embedding dimensions. (3) bge-reranker-base Xiao et al. (2023): a powerful cross-encoder architecture reranker proposed by Beijing Academy of Artificial Intelligence, with 278M parameters.
3.4 Implementation Details
We implement RADIO on FlashRAG Jin et al. (2024), a Python library for efficient RAG research. In the RAG pipeline, we take e5-base-v2 Wang et al. (2022) as the retriever, and Meta-Llama-3.1-8B-Instruct Touvron et al. (2023) as the generator. We sample 20,000 instances from NQ and TriviaQA to construct the fine-tuning dataset and fine-tune rerankers separately. For document sampling, we set the shift $n$ in Top-k shifted by N method as 3, and sample 6 negative samples from the subsequent documents. To ensure a fair comparison, the sampling index is fixed and remains unchanged across methods. In the RAG pipeline, we set the number of documents selected by retriever and reranker (i.e., $k_{1}$ and $k_{2}$) as 20 and 5. For fine-tuning the reranker, we tune the training epochs from 1 to 5 and the integration hyperparameter $\alpha$ from 0.0 to 1.0. We use Adam Kingma (2014) optimizer with a learning rate 6e-5 and a weight decay of 0.01. The prompts we used in experiments are given in Appendix A.4.
3.5 Main Results
Table 3: Transferability analysis across generators. EM is reported as the metric for MMLU dataset.
Generator Method NQ MMLU
EM F1 Humanities Social STEM Other ALL
Llama3.1-8b-instruct Base 0.3371 0.4603 0.4089 0.6867 0.5147 0.6650 0.5502
RADIO (ours) 0.3665 0.4917 0.4172 0.7013 0.5080 0.6717 0.5562
qwen2.5-14b-instruct Base 0.3310 0.4484 0.5439 0.8200 0.7206 0.7541 0.6906
RADIO (ours) 0.3518 0.4753 0.5598 0.8229 0.7250 0.7631 0.6995
gpt4o-mini Base 0.3607 0.4880 0.6485 0.8362 0.6784 0.8005 0.7300
RADIO (ours) 0.3742 0.5086 0.6548 0.8372 0.6768 0.8010 0.7321
3.5.1 Open-domain QA
To evaluate the effectiveness of RADIO and its transferability across different rerankers, we conduct experiments on the NQ and TriviaQA datasets. The results are presented in Table 1. From these results, we can draw the following conclusions:
  • Compared to the base method, most experimental settings achieve better results, demonstrating the necessity of preference alignment within the RAG pipeline.
  • Compared to other baseline methods, RADIO consistently achieves superior performance across all datasets and reranker backbone configurations. The results validate the effectiveness of using rationales as signals for preference alignment.
  • On TriviaQA, methods such as RRR and REPLUG show performance declines relative to the base method when using rerankers gte-large and bge-reranker-base. This indicates that these methods are sensitive and have limited applicability. In contrast, RADIO demonstrates robust adaptability to different rerankers, achieving significant performance improvements across all three rerankers.
  • As the reranker becomes larger or more powerful (e.g., progressing from gte-base to gte-large and further to bge-reranker-base), the performance ranking of models fine-tuned with RADIO aligns with the reranker’s inherent capabilities. This suggests that RADIO’s performance gains are sustainable and scalable with stronger rerankers, providing an avenue to further explore the upper performance limits of RAG pipelines.
3.5.2 MMLU
We also conduct experiments on MMLU. Since MMLU is a multiple-choice dataset, we report the EM metric Ma et al. (2023). Additionally, following previous work Yu et al. (2023), we fine-tune the reranker using open-domain QA as the source task and evaluate its performance on the MMLU dataset. The results are shown in Table 2 and we can draw the following conclusions:
  • From the metrics corresponding to the ALL category, RADIO demonstrates consistent improvements over the base method. This highlights the effectiveness and transferability of RADIO, as it successfully adapts to multi-choice question tasks even when fine-tuned on the Open-domain QA tasks.
  • Analyzing the results by question category, RADIO shows more significant improvements over the base method in the Humanities and Social Sciences categories, with average gains of 2.74% and 1.61%, respectively. However, it exhibits a slight negative effect in the STEM category. This may be due to the fine-tuning datasets (NQ and TriviaQA), which are Open-domain QA datasets with distributions more similar to humanities and social sciences but markedly different from STEM subjects.
  • Compared to other baseline methods, RADIO achieves top performance in the vast majority of metrics, demonstrating its superiority and state-of-the-art capability.
3.6 Transferability Analysis across Generators
We conduct experiments on two datasets of different tasks, NQ and MMLU, using three different generators (Llama3.1-8b-instruct Touvron et al. (2023), qwen2.5-14b-instruct Yang et al. (2024), and gpt4o-mini Achiam et al. (2023)) to validate the transferability of our method, as shown in Table 3. We can find: (1) RADIO maintains its effectiveness across different generators, consistently enhancing the performance of the original RAG pipeline. This demonstrates RADIO’s strong transferability with various generators. (2) Comparing different generators reveals that RADIO’s performance gains are more pronounced with smaller, less capable generators. Specifically, when the generators are Llama3.1, Qwen2.5, and GPT4o-mini, RADIO achieves EM improvements of 8.72%, 6.28%, and 3.74%, respectively, and F1 improvements of 6.82%, 6.00%, and 4.22%. This is because as the generator’s capability increases and approaches the upper performance limits of the RAG pipeline, further enhancing the pipeline becomes increasingly challenging, resulting in a smaller improvement.
3.7 Ablation Study
Table 4: Ablation study.
Dataset Metrics w/o ALL w/o Retrieval RADIO
NQ EM 0.3371 0.3587 0.3665
F1 0.4603 0.4858 0.4917
MMLU Humanities (EM) 0.4089 0.4168 0.4172
Social (EM) 0.6867 0.6981 0.7013
STEM (EM) 0.5147 0.5109 0.508
Other (EM) 0.665 0.6666 0.6717
ALL (EM) 0.5502 0.5548 0.5562
To explore the specific impact of rationale and retrieval score, we design the following variants: (1) w/o ALL: Base reranker without fine-tuning. Do not introduce rationale or retrieval score. (2) w/o Retrieval: Ranking documents and fine-tuning reranker only based on the rationale scores. (3) RADIO: Fine-tuning reranker based on both rationale and retrieval scores.
Table 4 shows the results of ablation study on NQ and MMLU, where we can derive the following findings: (1) Both the rationale score and retrieval score contribute positively to RADIO’s performance, with the rationale score demonstrating a stronger positive impact compared to the retrieval score. (2) RADIO outperforms both w/o ALL and w/o Retrieval, while w/o Retrieval surpasses w/o ALL. This indicates that the rationale score and retrieval score are not conflicting but rather complementary. Their integration provides a more robust signal for document ranking, which effectively aids in fine-tuning the reranker. (3) In the STEM category of the MMLU dataset, w/o ALL outperforms both w/o Retrieval and RADIO. This could be attributed to the fact that the training dataset (NQ) contains questions with distributions more similar to humanities and social sciences, leading to trends in the Humanities, Social Sciences, and Other categories that differ from STEM category.
3.8 Hyperparameter Analysis
Figure 2: Hyperparameter analysis on NQ and MMLU.
Figure 2 visualizes the performance of RADIO across different integration coefficients $\alpha$ on NQ and MMLU. The x-axis represents the integration coefficients $\alpha$ and the y-axis represents the evaluation metrics EM (red) or F1 (blue). For the MMLU dataset, we present results for two representative categories: Humanities and STEM. The trends for other categories align with those observed in Humanities. Complete experimental results are provided in Appendix A.3 for reference. From the figure, we can draw the following conclusions: (1) As $\alpha$ increases, the metrics on the NQ dataset and most categories of the MMLU dataset exhibit a trend of first rising and then falling, with the optimal range for $\alpha$ being between 0.3 and 0.7. This demonstrates the complementary nature of rationale scores and retrieval scores, which together form an optimal signal for preference alignment. (2) When $\alpha=0$, the RAG performance is suboptimal because document scoring relies entirely on retrieval scores, focusing solely on query-document relevance while ignoring whether the document supports the generator in answering the query. Conversely, when $\alpha=1$, the performance is still not optimal, as it completely disregards retrieval relevance, leading to a mismatch between the fine-tuning dataset and the training data, which negatively affects model performance. (3) The trends for STEM differ from those of other MMLU categories, showing an opposite pattern. This is likely due to the significant distributional differences between STEM and Humanities/Social Sciences, resulting in a "seesaw effect" as observed in the figure. This phenomenon is also reflected in REPLUG Shi et al. (2023), where the improvement in the STEM category is weaker compared to other categories.
3.9 Case Study
Figure 3: Case study on NQ dataset.
To intuitively illustrate the effectiveness of RADIO, we select examples from the NQ dataset to compare the documents reranked by RADIO with those reranked by a novel baseline, DPA-RAG, as well as their responses. In Figure 3 Example 1, the query asks, "Which state is the richest state in Nigeria?" RADIO successfully ranks information about Lagos State’s economic and financial status, which relates to the correct answer, among the top-3 documents. In contrast, DPA-RAG fails to identify documents relevant to answering the query, and cannot provide a correct response. In Figure 3 Example 2, the query is, "Who is the highest-selling R&B artist of all time?" RADIO prioritizes documents containing information about the correct answer, Michael Jackson, and effectively highlights key terms such as "R&B" and "best-selling." However, DPA-RAG misinterprets the query’s constraints, retrieving documents that either overlook the R&B artist specification or fail to consider the time span, resulting in an incorrect response. These examples demonstrate that RADIO enhances the RAG pipeline by providing more efficient and accurate reranking. It selects contextually appropriate documents, enabling the generator to infer the correct answers effectively.
4 Related Work
4.1 Retrieval-Augmented Generation
Large language models (LLMs) have demonstrated groundbreaking performance across numerous NLP tasks but still face challenges such as hallucination and outdated knowledge Gao et al. (2023). To address these issues, retrieval-augmented generation (RAG) has been introduced Fan et al. (2024). RAG retrieves relevant information from external knowledge bases and incorporates it as contextual input to the generator (LLM), enhancing the accuracy and reliability of the generated responses. The typical RAG pipeline can be divided into several key components: query rewriter, retriever, reranker, and generator. The query rewriter Wang et al. (2023) modifies and expands the original query to improve retrieval recall, ensuring more relevant documents are retrieved. The retriever Chen et al. (2024) fetches relevant documents based on the query. Dense retrievers generally outperform sparse retrievers in this step. To integrate contextual information more effectively and identify documents more relevant to the query, rerankers Moreira et al. (2024a) with larger models and greater complexity are introduced to reorder the retrieved documents compared to retrievers. Finally, the generator—usually a powerful LLM such as GPT-4 Achiam et al. (2023) or Llama Touvron et al. (2023)—uses the query and the top-$k$ documents from the reranker to generate the final response.
4.2 Preference Alignment
To further improve LLMs, preference alignment is often performed after the initial pretraining phase Jiang et al. (2024). Approaches such RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2024) are proposed to align the output of LLMs more closely with human preferences. DPO transforms tasks into classification problems, achieving high computational efficiency and strong performance. In the context of RAG, several works can be transformed to address the preference alignment challenge across RAG components. REPLUG Shi et al. (2023) improves RAG pipelines involving black-box LLMs by using the probability of the LLM generating the correct answer as a signal to determine document importance. Similarly, RRR Ma et al. (2023) uses metrics based on the quality of the LLM’s generated response as a signal to evaluate a document’s utility. On the other hand, ARL2 Zhang et al. (2024) prompts LLMs to generate self-guided relevance labels for fine-tuning retriever, and DPA-RAG Dong et al. (2024) introduces a bidirectional alignment strategy to mitigate preference inconsistencies in RAG pipelines. In this work, we focus on optimizing the reranker within RAG. Our goal is to enable the reranker to effectively identify supportive documents well-suited for the generator, whether black-box or open-source, facilitating the production of accurate outputs. In addition, BGM Ke et al. (2024) trains a bridge model between retriever and LLMs to transform the retrieved information into the format LLMs prefer, which can be used with RADIO in combination.
5 Conclusion
In this paper, we propose a novel and practical preference alignment framework, RADIO, with rationale distillation in retrieval-augmented generation. First, we introduce a rationale extraction method to extract the rationales necessary for answering queries with LLMs. Next, a rationale-based alignment is proposed to rerank documents based on extracted rationales and fine-tune rerankers. Extensive experiments are conducted to verify the effectiveness of our proposed method compared to other state-of-the-art baselines.
6 Limitations
First, compared to other methods, our approach RADIO requires additional time in the rationale extraction stage to generate rationales. Since different samples are independent of one another, we can reduce generation time by employing parallel processing to mitigate this issue. Secondly, the MMLU experimental results reveal that the composition of fine-tuning datasets can affect RADIO’s effectiveness. This issue can be addressed by designing task-specific fine-tuning datasets for different downstream tasks or large-scale general fine-tuning datasets.
Appendix A Appendix
Figure 4: More results of hyperparameter analysis on MMLU.
A.1 Dataset Desctiptions
The detailed descriptions of baselines are given as follows:
  • Natural Questions (NQ): NQ contains real user questions compiled from Google search, with corresponding answers identified from Wikipedia by human annotators.
  • TriviaQA: TriviaQA dataset comprises trivia questions paired with answer annotations and supporting evidence documents, such as web pages and Wikipedia articles. It is designed to assess a model’s ability to retrieve and comprehend textual evidence for open-domain question answering.
  • Massive Multitask Language Understanding (MMLU): MMLU is a comprehensive evaluation dataset comprising 57 categories of questions, which are grouped into four broad domains: Humanities, Social Sciences, STEM, and Other. In this paper, we report evaluation metrics based on these categories.
A.2 Baselines
Following is the introduction of baselines:
  • Base Xiao et al. (2023): The reranker model is used off-the-shelf without any fine-tuning.
  • Atlas Izacard et al. (2023): A pretrained retrieval-augmented language model designed for knowledge intensive task. We choose the EMDR2 Singh et al. (2021) as the reward to rerank documents and fine-tune rerankers.
  • REPLUG Shi et al. (2023): REPLUG seeks to fine-tune the retriever to enhance RAG pipelines that include black-box LLMs. It achieves this by using the query and document as contextual inputs and leveraging the probability of the LLM generating the correct answer as the importance score. This idea is also reflected in the PDist method in Atlas Izacard et al. (2023).
  • Trainable rewrite-retrieve-read (RRR) Ma et al. (2023): RRR optimizes the query rewriter using the evaluation metrics of the final RAG output as a reward, which is used to fine-tune rerankers in our pipeline, enhancing the overall effectiveness of RAG.
  • ARL2 Zhang et al. (2024): ARL2 introduces a method to use LLMs as supervisor to generate self-guided relevance labels for fine-tuning retriever.
  • DPA-RAG Dong et al. (2024): DPA-RAG proposes a knowledge preference pipeline to dual-align rerankers and generators in RAG. It combines document importance from the LLM’s perspective with the importance determined during the retrieval stage.
A.3 More Results on Hyperparameter Analysis
Table 5: Prompts for Open-domain datasets.
System Prompt: Answer the question based on
the given document. Only give me the answer
and do not output any other words. The
following are given documents.
{reference}
User Prompt:
Question: {question}
Answer:
Table 6: Prompts for Multi-choice datasets.
System Prompt: Answer the question based on
the given document. Only give me the option
(A/B/C/D) and do not output any other words.
The following are given documents.
{reference}
User Prompt:
Question: {question}
Answer:
Figure 4 illustrates the trend of RADIO’s performance across various MMLU categories as the integration coefficient $\alpha$ increases. For Humanities, Social Sciences, and Other, the trends are consistent: as $\alpha$ increases, the performance metrics first improve and then decline. However, the STEM category shows a unique pattern, with metrics initially decreasing as $\alpha$ grows, followed by an improvement. This divergence may be attributed to the fine-tuning dataset (NQ), which shares a closer distribution with Humanities, Social Sciences, and Other categories, while differing significantly from the STEM category.
A.4 Prompts
In this section, we detail the prompts we used in the experiments. For Open-domain QA datasets NQ and TriviaQA, the prompts are as shown in Table 5. For Multi-choice dataset MMLU, the prompts are shown in Table 6.